# Bayesian evidence for two peaks in the sound speed in cold dense QCD

### v3

@@ Appendix D Illustrations of favored and disfavored scenarios
Note that eq. 13 does not explicitly require $C_{s,\mathrm{max}}^{\mathrm{NS}}$ to be greater than $1/3$, though astrophysics strongly suggests the violation of conformal limit inside NSs. Moreover, while $C_{s}(n_{B})$ is allowed to have a finite number of discontinuities, a fact implicit in the central results  eq. 13 and in NS models in  appendix B (via very short length scales $\tilde{l}$), for aesthetic reasons I only show smooth curves here. Finally, at low densities the possibility of a peak in $C_{s}$ in the ${\chi\text{EFT}}$-based outer core or in the crust is not counted towards the number of peaks in this work.

### v4

@@ Appendix D Illustrations of favored and disfavored scenarios
Note that eq. 13 does not explicitly require $C_{s,\mathrm{max}}^{\mathrm{NS}}$ to be greater than $1/3$, though astrophysics strongly suggests violation of conformal limit inside NSs. Moreover, while $C_{s}(n_{B})$ is allowed to have a finite number of discontinuities, a fact implicit in the central results  eq. 13 and in NS models in  appendix B (via very short length scales $\tilde{l}$), for aesthetic reasons I only show smooth curves here. Finally, at low densities the possibility of a peak in $C_{s}$ in the ${\chi\text{EFT}}$-based outer core or in the crust is not counted towards the number of peaks in this work.